OpenStudy (anonymous):

A charter airline finds that on its Saturday flights from Philadelphia to London, all 120 seats will be sold if the ticket price is $200. However, for each $3 increase in ticket price, the number of seats sold decreases by one. (a) Find a formula for the number of seats sold if the ticket price is P dollars.

Directrix (directrix):

Suggested solution by @Theo5891 This formula solves for the number of tickets sold based on the price per ticket. that formula is: T = 120 - (P - 200) / 3 Testing the formula: If the price per ticket is 200, the number of tickets sold is 120. If the price per ticket is 203, the number of tickets sold is 120 - (3/3) = 120 - 1 = 119. If the price per ticket is 206, the number of tickets sold is 120 - (6/3) = 120 - 2 = 118.